# PETSc/TAO Developments for Early Exascale Systems

### v1

@@ 6.2 Improve asynchronicity with PetscDeviceContext and managed memory
Conceptually, device context is an abstraction for vendor GPU streams. It is comprised of the device over which it presides and the stream that it manages, but also bears other responsibilities. We provide APIs for one to synchronize a device context, or make one context wait for another. With vendor programming models, the inter-stream data dependence is usually done via event objects (e.g., cudaEvent_t) where a data producer records an event on a stream and a data consumer waits for the event on another stream. But using event objects requires users to store or pass the objects around, easily messing up the library call chain. Instead, we hide the complexity within PETSc device contexts and only expose high-level context control APIs. We provide memory registration APIs to assign memory regions a PETSc object ID. With that, we can connect PETSc GPU vector/matrix objects and memory regions with device contexts. It is done via a pair of Begin/End marking APIs that take arguments of a device context, a PETSc object ID, and an access mode as below:
@@
They notify the runtime the PETSc object (or memory region) in the enclosed scope is accessed read-only, write-only or read-write. Once an object has been marked, subsequent accesses to it are ordered according to its data dependencies. Operations using the same device context are ordered and executed as they appear in the source code. On the other hand, operations using separate device contexts which access a common object are strongly write-ordered. That is, operations are serialized by the order in which they are executed. Formally, given two operations A and B and an object, if A writes and B reads the object, and A appears before B in source code, then B shall not begin before A completes. The rule also holds for write-after-write or write-after-read. To enforce write-ordering, PETSc will internally create event objects for all input and output objects of an operation. The latest write event or all the read events since the last write are recorded after the operation has launched; they are used as semaphores for subsequent launches. If an object has previously recorded events, and the new operation conflicts with the previous operation (by rules defined above), then PETSc will ensure that the current operation does not begin until those events have completed. It is important to note that waiting for the previous operation is done asynchronously on the device whenever possible.
PETSc managed memory is used to facilitate managing a small collection of scalar variables. These may be needed to mutate or scale larger objects or, in linear solvers, for example, to determine algorithm convergence. Often, they are produced by some routine, trivially modified or inspected, and then passed directly on to a consumer. In the subsection above, we used raw device pointers for scalar variables. Managed memory models use a std::future for an opaque, host/device mirrored dual array of values. Based on that, subclass ManagedReal is used to manipulate real-valued scalar variables on device or host. ManagedReal has the ability to symbolically represent expressions and convert them into the corresponding GPU or CPU kernels by using C++ expression templates. These symbolic expressions are then executed on the GPU or CPU, depending on runtime options. An example asynchronously normalizing a vector is shown in Listing 5.
@@
7 Algebraic multigrid on GPUs
Performance portability is demonstrated by a scaling study with PETSc’s built-in algebraic multigrid (AMG) solver, PCGAMG, using the PETSc/Kokkos backend over CUDA and HIP. PCGAMG uses the PETSc multigrid PCMG framework and can thus take advantage of optimized backend operations. This ability to abstract the AMG algorithm with standard sparse linear algebra has facilitated its widespread use in the PETSc and wider computational science communities. PETSc’s built-in functionality for finite elements is used to discretize the Laplacian operator with second-order elements. Each MPI process has a logical cube of hexahedral cells, with 64 such processes per node (e.g., 16 MPI tasks per A100). Increasingly larger grids are generated by uniform refinements.
This data was generated with a test harness111src/snes/tests/ex13.c that uses a two-level partitioning of Cartesian grids in DM, first to nodes and then to MPI tasks on each node. These experiments were run with 64 processes per node.
8 Batched linear solvers
Batching is a technique for exposing PE-level parallelism in algorithms that previously ran on entire processes or multiple threads within a single MPI process. Kinetic discretizations of magnetized plasmas, for example, advance the Vlasov-Maxwell system, which is then followed by a fully implicit time advance of a collision operator. These collision advances are independent at each spatial point and are well suited to batch processing. The full implicit time integrator in our Landau operator (Section 10) requires linear solves that can effectively run these many small systems on GPUs. PETSc has developed batched version of two Krylov methods, TFQMR and BiCG, with diagonal preconditioning. Batched iterative solvers of this form have also been deployed in the Kokkos-Kernels and Ginkgo libraries (Liegeois et al., 2023; Anzt et al., 2022). Figure 9 shows the solves per second, as a function of batch size, of the batched and “ensemble” solvers. Ensemble solvers simply stack the individual systems in one large system and use standard PETSc solvers. A 3X performance increase is observed with batching. For details see Adams et al. (2024).
@@ 8 Batched linear solvers
9 Dense reformulations
Reformulating algorithms to replace level-1 BLAS operations with level-2 or level-3 operations generally improves performance, but dense reformulations are more challenging to formulate for flexible algorithms that use callbacks that cannot be vectorized. In PETSc/TAO we have developed a dense reformulation of the limited-memory BFGS method (L-BFGS, Liu and Nocedal (1989)), one of the most popular optimization algorithms, that improves performance while maintaining the flexibility of the original implementation.
@@ 9 Dense reformulations
The history size $m$ is fixed, and the method requires the storage of history vectors $S_{k}=[s_{k-m}|\cdots|s_{k-1}]$ and $Y_{k}=[y_{k-m}|\cdots|y_{k-1}]$ and the precomputation of the dot products $d_{j}:=s_{j}^{T}y_{j}$. In the implementation of the recursive algorithm in PETSc/TAO, $H_{0}^{(k)}g$ is computed as a callback to an arbitrary linear operator. This operator is not only arbitrary but may change from iteration to iteration, which allows this implementation of L-BFGS to support variable-metric methods (Davidon, 1991). For many problems, variable-metric L-BFGS has been shown to converge faster than L-BFGS with a constant base operator $H_{0}^{(k)}\equiv H_{0}$ (Dener and Munson, 2019). The recursive algorithm is work optimal, but it includes $2m$ synchronizations (one for each dot product) and, as it is based on level-1 BLAS, has little temporal locality that takes advantage of the cache hierarchy.
@@
For small problem sizes ($n\leq 10^{3}$ on the CPUs, $n\leq 10^{5}$ on the GPUs), when the latency of synchronization is the dominant cost, the intermediate dense method exhibits runtimes that are almost independent of $m$, unlike the recursive algorithm with its $2m$ synchronizations. For the largest problems, when a single vector does not fit into the last level of cache, a rank-1 update cannot cache the vector that is being updated, so the $2m$ rank-1 updates of the recursive method will require almost twice the memory traffic of two rank-$m$ updates of the intermediate dense method. We see this in the effective bandwidth for the largest problem sizes on both CPUs and GPUs, and the same reasoning explains why the intermediate dense method achieves almost twice the bandwidth of the recursive dense method when all of the problem data fits into the last level of cache for CPU computations, such as in Figure 10 (top) when $n=10^{5}$ and $m=50$. This data shows that the intermediate dense L-BFGS formulation in PETSc/TAO exhibits almost uniformly superior performance to the recursive formulation while retaining all of the flexibility of the recursive formulation in supporting variable-metric methods. Further dense reformulations of iterative methods that retain flexibility are planned as future improvements to PETSc/TAO.
10 Fully GPU enabled time evolution of Landau collision integral
PETSc provides GPU support for the entire PDE solver stack, from time integrators to nonlinear solvers, batched linear solvers, and COO matrix assembly. A fully GPU enabled Landau collision time advance is implemented entirely in PETSc with “mini-app” drivers as a PETSc example (Hirvijoki and Adams, 2017; Adams et al., 2017, 2022b, 2024). The structure of this code is shown in Figure 11.
@@ 10 Fully GPU enabled time evolution of Landau collision integral
This method is an example of building an entire HPC PDE solver in PETSc, with a small driver code222 src/ts/utils/dmplexlandau/tutorials/ex2.c in the PETSc repository mimicking an application, with verification tests, used for performance experiments, and a specific PDE operator (“Landau”) that would, in general, be in user code.
@@ 10.2 Validation of Landau collision integral
A common anisotropic plasma verification test has recently been published that initializes a two species plasma (electrons and deuterium) with (4) different temperatures for the parallel and perpendicular temperature of each species that is allowed to evolve toward equilibrium. A driver code runs many of these problems simultaneously to mimic its use in an application where 1,000s of spatial vertices would be processed simultaneously. This problem uses PETc’s adaptive time stepping with over 14,000 time steps to achieve near full thermalization, using new high-order simplex finite elements and verification with analytical results of thermalization rates (Adams et al., 2024). Figure 13 shows a temperature history for the $P2$ case.
@@ 11 Conclusion and Vision for the Future
We have summarized some of the challenges encountered and the advances made toward providing performance-portable support for GPUs in PETSc over the course of ECP. This effort has required development ranging across the different levels of the PETSc software stack. As accelerated GPU computation continues to shape the HPC field, we will continue to improve PETSc’s capabilities for this, adding new features (e.g., support for the block CSR matrix format on GPUs), and offloading more computations. We hope to consolidate PETSc GPU backends to simplify implementation and maintenance, and we will continue to collaborate with developers of PETSc external libraries to smooth the interface for GPU data passing between libraries. An important direction that PETSc developers are currently pursuing is adding robust support for machine learning computations in PETSc to support a broad range of applications, which demands high GPU performance.
Maintaining a stable interface is important for a numerical library like PETSc/TAO, but our experience has shown the tension between that goal and the goal of portable performance for pre- and early exascale machines. In some cases, as in the dense reformulation of L-BFGS, additional parallelism can be found within existing interfaces that is portable across systems; in other cases, as in the introduction of the ManagedMemory type, something that is technically a change to the library’s interface finds additional parallelism within existing programming patterns used by applications. First in introducing PetscSF and now in MatSetValuesCOO(), PETSc has had success with new interfaces that (a) are declarative, (b) expose as much of the parallelism within the desired computation as possible, and (c) are independent of the computational resources used to execute the operation and the memory resources holding the arguments. Once introduced, interfaces with these properties appear to become stable features of the library, and applications that adopt them should see benefits. In the future, continued evolution of PETSc/TAO along these lines will help applications achieve portable performance for larger and larger percentages of their workflows, both on exascale systems and those that will follow.
@@
The authors were supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under Contract DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357, resources of the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, and resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. The work was partially done on a pre-production supercomputer with early versions of the Aurora software development kit. The authors were partially supported by the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration. MGK was partially supported by NSF CCSI award 1931524.

### v2

@@ 6.2 Improve solver asynchronicity with PetscDeviceContext and managed memory
Conceptually, a PETSc device context is an abstraction for vendor GPU streams (not to be confused with the CUcontext type in CUDA driver, which is an encapsulation for all resources and actions performed within the CUDA driver API). It encodes the device ID over which it presides and the stream it manages, but also bears other responsibilities. We provide APIs for one to create, synchronize, or fork and join device contexts. With vendor programming models, the inter-stream data dependence is usually done via event objects (e.g., cudaEvent_t) where a data producer records an event on a stream and a data consumer waits for the event on another stream. But using event objects requires users to store or pass the objects around, rendering the library call chain error-prone. We now hide the complexity within PETSc device contexts and expose higher level context control APIs. In addition, we provide memory registration APIs to assign memory regions a PETSc object ID. With that, we can connect PETSc GPU Vec/Mat objects and memory regions with device contexts. It is done via a pair of Begin/End marking APIs that take arguments of a device context, a PETSc object ID, and an access mode as below:
@@
They notify the runtime the PETSc object (or memory region) in the enclosed scope is accessed read-only, write-only or read-write. Once an object has been marked, subsequent operations on it are ordered to respect data dependency, while independent operations might be executed in parallel with multiple contexts. To enforce data dependency, PETSc will internally create event objects for all input and output objects of an operation. The latest write event or all the read events since the last write are recorded after the operation has been launched; they are used as semaphores for subsequent launches. If an object has previously recorded events, and the new operation conflicts with a previous one, then PETSc will ensure that the current operation does not begin until those events have completed. It is important to note that waiting for the previous operation is done asynchronously on the device whenever possible.
In PETSc, we often need scalars to mutate or scale larger objects or, in linear solvers, to determine algorithm convergence. These scalars might be output by a kernel, trivially modified or inspected, and then passed directly on to another kernel. In the subsection above, we used raw device pointers to scalar variables, resulting in non-portable code. To overcome this, we designed Petsc::ManagedMemory, which uses a std::future for an opaque, host/device mirrored dual array of values. For example, its subclass ManagedReal is used to manipulate real-valued scalar variables on device or host. ManagedReal has the ability to symbolically represent expressions, convert them into the corresponding CPU or GPU kernels by using C++ expression templates, and then execute them on host or device, depending on runtime options. An example asynchronously normalizing a vector is shown in Listing 5.
@@
7 Batched linear solvers on GPUs
Conceptually related (and complementary to) the use of asychronicity in linear solvers is the use of batching. Batching is a technique for exposing PE-level parallelism in algorithms that previously ran on entire processes or multiple threads within a single MPI process. Kinetic discretizations of magnetized plasmas, for example, advance the Vlasov-Maxwell system, which is then followed by a fully implicit time advance of a collision operator. These collision advances are independent at each spatial point and are well suited to batch processing. The full implicit time integrator in our Landau operator (Section 10) requires linear solves that can effectively run these many small systems on GPUs. PETSc has developed batched version of two Krylov methods, TFQMR and BiCG, with diagonal preconditioning. Batched iterative solvers of this form have also been deployed in the Kokkos-Kernels and Ginkgo libraries (Liegeois et al., 2023; Anzt et al., 2022). Figure 7 shows the solves per second, as a function of batch size, of the batched and “ensemble” solvers. Ensemble solvers simply stack the individual systems in one large system and use standard PETSc solvers. A 3X performance increase is observed with batching. For details see Adams et al. (2024).
@@ 7 Batched linear solvers on GPUs
8 Dense reformulations of algorithms
Reformulating algorithms to replace level-1 BLAS operations with level-2 or level-3 operations generally improves performance, but dense reformulations are more challenging to formulate for flexible algorithms that use callbacks that cannot be vectorized. In PETSc we have developed a dense reformulation of the limited-memory BFGS method (L-BFGS, Liu and Nocedal (1989)), one of the most popular optimization algorithms, that improves performance while maintaining the flexibility of the original implementation.
@@ 8 Dense reformulations of algorithms
The history size $m$ is fixed, and the method requires the storage of history vectors $S_{k}=[s_{k-m}|\cdots|s_{k-1}]$ and $Y_{k}=[y_{k-m}|\cdots|y_{k-1}]$ and the precomputation of the dot products $d_{j}:=s_{j}^{T}y_{j}$. In the implementation of the recursive algorithm in PETSc, $H_{0}^{(k)}g$ is computed as a callback to an arbitrary linear operator. This operator is not only arbitrary but may change from iteration to iteration, which allows this implementation of L-BFGS to support variable-metric methods (Davidon, 1991). For many problems, variable-metric L-BFGS has been shown to converge faster than L-BFGS with a constant base operator $H_{0}^{(k)}\equiv H_{0}$ (Dener and Munson, 2019). The recursive algorithm is work optimal, but it includes $2m$ synchronizations (one for each dot product) and, as it is based on level-1 BLAS, has little temporal locality that takes advantage of the cache hierarchy.
@@
For small problem sizes ($n\leq 10^{3}$ on the CPUs, $n\leq 10^{5}$ on the GPUs), when the latency of synchronization is the dominant cost, the intermediate dense method exhibits runtimes that are almost independent of $m$, unlike the recursive algorithm with its $2m$ synchronizations. For the largest problems, when a single vector does not fit into the last level of cache, a rank-1 update cannot cache the vector that is being updated, so the $2m$ rank-1 updates of the recursive method will require almost twice the memory traffic of two rank-$m$ updates of the intermediate dense method. We see this in the effective bandwidth for the largest problem sizes on both CPUs and GPUs, and the same reasoning explains why the intermediate dense method achieves almost twice the bandwidth of the recursive dense method when all of the problem data fits into the last level of cache for CPU computations, such as in Figure 8 (top) when $n=10^{5}$ and $m=50$. This data shows that the intermediate dense L-BFGS formulation in PETSc exhibits almost uniformly superior performance to the recursive formulation while retaining all of the flexibility of the recursive formulation in supporting variable-metric methods. We plan to study the benefits of these improved performances for training deep neural networks (Zampini et al., 2024). Further dense reformulations of iterative methods that retain flexibility are planned as future improvements to PETSc.
9 Mini-application study 1: Solving 3D Laplacian with algebraic multigrid
In this section we demonstrate the performance portability of the PETSc/Kokkos backend over CUDA and HIP with a scaling study of the PETSc’s built-in algebraic multigrid (AMG) solver PCGAMG. PCGAMG uses the PETSc multigrid PCMG framework and can thus take advantage of optimized backend operations. This ability to abstract the AMG algorithm with standard sparse linear algebra has facilitated its widespread use in the PETSc and wider computational science communities. PETSc’s built-in functionality for finite elements is used to discretize the Laplacian operator with second-order elements. Each MPI process has a logical cube of hexahedral cells, with 64 such processes per node (e.g., 16 MPI tasks per A100). Increasingly larger grids are generated by uniform refinements.
This data was generated with a test harness111petsc/src/snes/tests/ex13.c that uses a two-level partitioning of Cartesian grids in DM, first to nodes and then to MPI tasks on each node. These experiments were run with 64 processes per node.
10 Mini-application study 2: Time evolution of Landau collision integral
Using PETSc’s GPU support for the entire PDE solver stack, from time integrators to nonlinear solvers, batched linear solvers, and COO matrix assembly, a fully GPU enabled Landau collision time advance is implemented with “mini-app” drivers as a PETSc example (Hirvijoki and Adams, 2017; Adams et al., 2017, 2022b, 2024). The structure of this code is shown in Figure 11.
@@ 10 Mini-application study 2: Time evolution of Landau collision integral
This method is an example of building an entire HPC PDE solver in PETSc, with a small driver code222 petsc/src/ts/utils/dmplexlandau/tutorials/ex2.c mimicking an application, with verification tests, used for performance experiments, and a specific PDE operator (“Landau”) that would, in general, be in user code.
@@ 10.2 Validation of Landau collision integral
A common anisotropic plasma verification test has recently been published that initializes a two-species plasma (electrons and deuterium) with (4) different temperatures for the parallel and perpendicular temperature of each species that are allowed to evolve toward equilibrium. A driver code runs many of these problems simultaneously to mimic its use in an application where 1,000s of spatial vertices would be processed simultaneously. This problem uses PETc’s adaptive time stepping with over 14,000 time steps to achieve near full thermalization, using new high-order simplex finite elements and verification with analytical results of thermalization rates (Adams et al., 2024). Figure 13 shows a temperature history for the $P2$ case.
@@ 11 Conclusion and Vision for the Future
We have summarized some of the challenges encountered and the advances made toward providing performance-portable support for GPUs in PETSc over the course of ECP. This effort has required development ranging across the different levels of the PETSc software stack. As accelerated GPU computation continues to shape the HPC field, we will continue to improve PETSc’s capabilities for this, adding new features (e.g., support for the block CSR matrix format on GPUs), and offloading more computations. We hope to consolidate PETSc GPU backends to simplify implementation and maintenance, and we will continue collaborating with developers of PETSc external libraries to smooth the interface for GPU data passing between libraries. An important direction that PETSc developers are currently pursuing is adding robust support for machine learning computations in PETSc to support a broad range of applications, which demands high GPU performance.
Maintaining a stable interface is important for a numerical library like PETSc, but our experience has shown the tension between that goal and the goal of portable performance for pre- and early exascale machines. In some cases, as in the dense reformulation of L-BFGS, additional parallelism can be found within existing interfaces that is portable across systems; in other cases, as in the introduction of the ManagedMemory type, something that is technically a change to the library’s interface finds additional parallelism within existing programming patterns used by applications. First in introducing PetscSF and now in MatSetValuesCOO(), PETSc has had success with new interfaces that (a) are declarative, (b) expose as much of the parallelism within the desired computation as possible, and (c) are independent of the computational resources used to execute the operation and the memory resources holding the arguments. Once introduced, interfaces with these properties appear to become stable features of the library, and applications that adopt them should see benefits. In the future, the continued evolution of PETSc along these lines will help applications achieve portable performance for larger and larger percentages of their workflows, both on exascale systems and those that will follow.
@@
The authors were supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research under Contract DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357, resources of the National Energy Research Scientific Computing Center, a Department of Energy Office of Science User Facility, and resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. The work was partially done on a pre-production supercomputer with early versions of the Aurora software development kit. The authors were partially supported by the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration. MGK was partially supported by NSF CSSI award 1931524. SZ acknowledges the support of the Extreme Computing Research Center, King Abdullah University of Science and Technology.